A total of 308 tickets were sold for the school play. The number of student tickets sold was three times the number of adult tickets sold . How many adult tickets were sold?
Answer by stanbon(75887)

A total of 308 tickets were sold for the school play. The number of student tickets sold was three times the number of adult tickets sold . How many adult tickets were sold?
----------------------------
Equations:
s + a = 308
s = 3a
------------------
Substitute for "s" and solve for "a":
3a+a = 308
4a = 308
a = 77 (# of adult tickets sold)
Since s = 3a, s = 3*77 = 231 (# Of students tickets sold)
==========================================================
Cheers,
Stan H.